OpenStudy (anonymous):

the wave length of a moving particle is inversely proportional to its momentum. if the momentum of the particle is decreased by 60%,find the percentage change in wave length! Note: Momentum is the product of mass and velocity someone can help me out?? :D

terenzreignz (terenzreignz):

two values, a and b are inversely proportional if and only if their product at any given time is constant. \[\large ab=k\]

OpenStudy (anonymous):

ooooo000ooo Thank You So Much,Terenz :D

terenzreignz (terenzreignz):

I doubt you needed help here at all :D

OpenStudy (anonymous):

haha,you're right.. this one,i can't answer it till you've told me

terenzreignz (terenzreignz):

Well, I'll answer it slowly... :D Let p be the momentum and let λ be the wavelength. Then \[\large p\lambda=k\]for some constant k.Now the momentum of the particle is decreased by 60%, so... replace p with p-0.6p And let μ be the new wavelength... \[\large (p-0.6p)\mu=p(1-0.6)\mu=0.4p\mu\] And this must still be equal to k \[\large 0.4p\mu=k\] k=pλ \[\large 0.4p\mu=p\lambda\]Cancelling p and dividing both sides by 0.4 \[\large \mu=2.5\lambda\] So the new wavelength μ is equal to 2.5λ, or λ + 150% of it. Case closed...

OpenStudy (anonymous):

aha! clearly enough now! Tx so much :D